OpenStudy (anonymous):

If y varies inversely as x, and y = 3 as x = -2, find y for the x-value of -1. WILL MEDAL

OpenStudy (nikato):

y varies inversely as x Is y=k/x And you know y=3 and x=-2 Plug them in and solve for k So 3=k/-2 Multiply both sides by -2 k=-6 So your equation is y= -6/x

OpenStudy (nikato):

And you know x=-1 And you're solving for y. So y= -6/-1 y=?

OpenStudy (anonymous):

6